Solve the given problems.

1. A population consists of the four numbers 1, 2, 4 and 5. List all the possible

samples of size n = 2

which can be drawn with replacement from the population.

Find the following:

a. Population mean

b. Population variance

c. Population standard deviation.

d. Mean of the sampling distribution of sample means

e. Variance of the sampling distribution of sample means

f. Standard deviation of the sampling distribution of sample means


1
Expert's answer
2022-05-06T12:13:56-0400

a. We have population values 1,2,4 and 5 population size N=4 and sample size n=2.

Mean of population "(\mu)" = 

"\dfrac{1+2+4+5}{4}=3"


b. Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}"


"=\dfrac{4+1+1+4}{4}=\dfrac{5}{2}"

c.

"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{5}{2}}\approx1.5811"

d. The number of possible samples which can be drawn with replacement is "2^4=16."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 &1,1 & 2/2 \\
 \hdashline
 2 & 1,2 & 3/2 \\
 \hdashline
 3 & 1,4 & 5/2 \\
 \hdashline
 4 & 1,5 & 6/2 \\
 \hdashline
 5 & 2,1 & 3/2 \\
 \hdashline
 6 & 2,2 & 4/2 \\
 \hdashline
 7 & 2,4 & 6/2 \\
 \hdashline
 8 & 2,5 & 7/2 \\
 \hdashline
 9 & 4,1 & 5/2 \\
 \hdashline
 10 & 4,2 & 6/2 \\
 \hdashline
 11 & 4,4 & 8/2 \\
 \hdashline
 12 & 4,5 & 9/2 \\
 \hdashline
 13 & 5,1 & 6/2 \\
 \hdashline
 14 & 5,2 & 7/2 \\
 \hdashline
 15 & 5,4 & 9/2 \\
 \hdashline
 16 & 5,5 & 10/2 \\
 \hdashline

\end{array}"



b.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) &\bar{X}^2 f(\bar{X})\\ \hline
 2/2 & 1/16 & 2/32 & 4/64\\
 \hdashline
 3/2 & 2/16 & 6/32 & 18/64\\
 \hdashline
 4/2 & 1/16 & 4/32 & 16/64\\
 \hdashline
 5/2 & 2/16 & 10/32 & 50/64\\
 \hdashline
 6/2 & 4/16 & 24/32 & 144/64\\
 \hdashline
7/2 & 2/16 & 14/32 & 98/64\\
 \hdashline
 8/2 & 1/16 & 8/32 & 64/64\\
 \hdashline
 9/2 & 2/16 & 18/32 & 162/64\\
 \hdashline
 10/2 & 1/16 & 10/32 & 100/64\\
 \hdashline
\end{array}"


Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=3=\mu"



e.The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{656}{64}-(3)^2=\dfrac{162}{40}=\dfrac{5}{4}= \dfrac{\sigma^2}{n}"

f.

"\sigma_{\bar{X}}=\sqrt{\dfrac{5}{4}}\approx1.1180"